# Perturbations of Toeplitz operators on vector-valued Hardy spaces

### v1

@@ Abstract.
In this article, we completely classify invariant subspaces of certain finite-rank perturbations of Toeplitz operators on vector-valued Hardy spaces. As a byproduct, we describe invariant subspaces, almost invariant subspaces, and nearly invariant subspaces associated with a class of Toeplitz operators in the vector-valued setting. Additionally, we investigate a generalization of the concept of nearly invariant subspaces for the finite defect case.
@@ 1. Introduction
Perturbations of Toeplitz operators on Hardy spaces have engendered a great deal of interest because–as is well known–they help us understand how small changes in these operators affect their properties, like spectrum or invertibility, which are crucial for solving equations in signal processing and control theory. They’re also key in studying stability in function spaces, where even tiny tweaks can impact convergence or approximation in applications like harmonic analysis or quantum mechanics. The basic ideas in this context are available in the volumes by Böttcher and Silbermann [BS] and Rosenblum and Rovnyak [RR1]. We also refer to the paper by Axler, Chang, and Sarason [AS].
@@
In this article, we explore and establish interesting properties and connections between perturbation theory, invariant subspaces, nearly invariant subspaces and almost invariant subspaces in the context of certain Toeplitz operators on vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. See Section 4 for the statements of our main results. It is worth mentioning that Hitt [Hit] introduced nearly invariant subspaces but called them weakly invariant. Subsequently, Sarason [Sar] rechristened them with the now universally accepted label nearly invariant subspaces and he introduced new ideas into their study by establishing a new proof via de Branges-Rovnyak spaces. A subspace $\mathcal{M}$ of the Hardy space $H^{2}(\mathbb{D})$ is called nearly invariant under the backward shift operator $T_{z}^{*}$ on $H^{2}(\mathbb{D})$ if $T_{z}^{*}(f)$ belongs to $\mathcal{M}$ whenever $f\in\mathcal{M}$ and vanishes at zero. Since then, numerous generalizations have been proposed in this direction. For further insights into nearly invariant subspaces, one can refer to [CCP], [CGP], [CD], [CDP1], [DS], [Era], [KLD], [LP], [LP1], [LP2], [Sar] and the references therein. We also refer to two recent related papers [CDP2] and [LP3].
@@
A transformative feature of nearly invariant subspaces of the backward shift operator on $H^{2}(\mathbb{D})$ observed and used by Sarason is their connection with a particular rank-one perturbation of the backward shift operator. Inspired by this masterful insight of Sarson, Das and Sarkar [DS] have recently investigated the invariant subspaces of certain Toeplitz operators, a generalization of nearly invariant subspaces, and almost invariant subspaces on $H^{2}(\mathbb{D})$ via a classification of invariant subspaces of some finite-rank perturbations of certain Toeplitz operators that they derive in the same paper. Our work in this paper is in alignment with the ideas from [DS], on which we shall elaborate more in the subsequent sections.
2. A Brief Preview
In the paper under consideration, we completely classify the invariant subspaces of certain finite-rank perturbations of Toeplitz operators on vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. Along the way, we explore rich and fascinating connections between perturbation theory, invariant subspaces, almost invariant subspaces, and nearly invariant subspaces within the framework of vector-valued Hardy spaces. Our methods are inspired by some ideas of Das and Sarkar proposed in [DS] for the scalar case; however, our proofs are self-contained and do not use any result from [DS], except for one observation (Lemma 5.5) that follows using some straightforward operator-theoretic arguments. Additionally, we study a generalization of the concept of nearly invariant subspaces for the finite defect case in the vector-valued setting.
The organization of the paper is as follows. In Section 3, we introduce the necessary notations, definitions, and terminologies that will be used throughout the paper. Section 4 presents the statements of our main results. Section 5 contains the proofs of Theorems 4.1 and 4.2 that classify invariant subspaces for certain finite-rank perturbations of Toeplitz operators on vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. Furthermore, it contains the proofs of three consequences of Theorems 4.1 and 4.2: (i) characterization of invariant subspaces of certain finite-rank perturbations of analytic Toeplitz operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$ (Theorem 4.3), (ii) characterization of almost invariant subspaces of certain Toeplitz operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$ (Theorem 4.5), (iii) characterization of nearly invariant subspaces associated with certain Blaschke-based operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$ (Theorem 4.7). Lastly, in Section 6, we give proofs of our results– Theorems 4.9, 4.10, and 6.3 –on nearly invariant subspaces with finite defect of vector-valued Hardy space. Moreover, Sections 5 and 6 also contain several significant remarks and crucial observations that highlight essential aspects of our findings.
@@ 3. Preliminaries
We shall use $\{E_{i}:1\leq i\leq n\}$ to denote the standard orthonormal basis of the Euclidean space $\mathbb{C}^{n}$. Note that each $E_{i}$ can be viewed as a function in $H^{2}(\mathbb{D},\mathbb{C}^{n})$ by identifying it with the constant function that is identically $E_{i}$. Furthermore, $\{E_{j}z^{m}:1\leq j\leq n,\ m\geq 0\}$ forms an orthonormal basis of the Hardy space $H^{2}(\mathbb{D},\mathbb{C}^{n}).$ For $n=1$, we denote $H^{2}(\mathbb{D},\mathbb{C}^{n})$ by $H^{2}(\mathbb{D})$.
@@
The latter identification allows one to view operators between these vector-valued Hardy spaces as row matrices. Suppose $\mathcal{G}$ is a bounded operator from $H^{2}(\mathbb{D},\mathbb{C}^{p})$ to $H^{2}(\mathbb{D},\mathbb{C}^{m})$ and $\mathcal{J}$ is a bounded operator from $H^{2}(\mathbb{D},\mathbb{C}^{r})$ to $H^{2}(\mathbb{D},\mathbb{C}^{m})$, then we shall use the notation $[\mathcal{G},\mathcal{J}]$ to denote the bounded operator from $H^{2}(\mathbb{D},\mathbb{C}^{p+r})$ to $H^{2}(\mathbb{D},\mathbb{C}^{m})$ defined by
@@
by a row matrix $[G_{1},\dots,G_{m}]$, and call it a matrix with $m$ columns $G_{1},\dots,G_{m}$.
@@
A bounded operator $T$ on a Hilbert space $\mathcal{H}$ is called a $C._{0}$ contraction if for each $h\in\mathcal{H}$, $\|T^{*n}h\|\longrightarrow 0$ as $n\longrightarrow\infty$. A $C._{0}$ isometry is called a pure isometry. We end this section by quoting an interesting result of Benhida and Timotin on $C._{0}$ contraction [BT] that is indispensable for our work in the following sections.
@@
[BT, Lemma 3.3] Suppose $T$ is a bounded operator on a Hilbert space $\mathcal{H}$. Let $T$ be a $C._{0}$ contraction such that the rank of $I-T^{*}T$ is finite, and let $\mathcal{M}\subset\mathcal{H}$ be a subspace of finite codimension. Then $TP_{\mathcal{M}}$ is a $C._{0}$ contraction, where $P_{\mathcal{M}}$ is the orthogonal projection of $\mathcal{H}$ onto $\mathcal{M}$.
@@ 4. Statements of the main results
The structural insights provided by Theorems 4.1 and 4.2 extend well beyond their immediate context and serve as a foundational tool for analyzing a variety of operator-theoretic phenomena on vector-valued Hardy spaces. We apply these to deduce the following three results.
@@
=$\mathcal{M}=\Big{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):(T^{*}_{F_{1}}F,\dots,T^{*}_{F_{p}}F,F)\in\mathcal{N}\Big{\}},$
@@
Suppose $B$ and $B^{{}^{\prime}}$ are two finite Blaschke products such that $B$ vanishes at zero, and $B$ divides $B^{{}^{\prime}}$ Let $\mathcal{M}$ be a non-zero nearly $T^{*}_{B,B^{{}^{\prime}}}$-invariant subspace of $H^{2}(\mathbb{D},\mathbb{C}^{m})$ with defect $1$, and let $J$ be a function of unit norm in the defect space.
@@
    =$\mathcal{M}=\Big{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=\mathcal{G}R+BhJ\ \ {\rm with}\ \ (R,h)\in\mathcal{K}\Big{\}},$
@@
    =$\mathcal{M}=\Big{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=BhJ\ \ {\rm with}\ \ h\in\mathcal{K}\Big{\}}.$
@@
    =$\mathcal{M}=\Big{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=\mathcal{G}R+BhJ\ \ {\rm for}\ (R,h)\in\mathcal{K}\Big{\}},$
@@
    =$\mathcal{M}=\Big{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=BhJ\ \ {\rm for}\ h\in\mathcal{K}\Big{\}},$
@@
5. Proofs of Theorems 4.1 & 4.2, and their Applications I, II & III
Before we give the proof of Theorem 4.1, first we state the following result from [DS]. It describes invariant subspaces of finite-rank perturbations of a Toeplitz operator $T^{*}_{\phi}$ on $H^{2}(\mathbb{D})$ with the associated symbol $\phi$ an inner function.
@@
[DS, Theorem 2.1] Suppose $\phi\in H^{2}(\mathbb{D})$ is an inner function that vanishes at zero; $\{u_{i}\}_{i=1}^{k}$ and $\{v_{i}\}_{i=1}^{k}$ are two orthonormal sets in $H^{2}(\mathbb{D})$. Let $\mathcal{M}$ be a non-zero subspace of $H^{2}(\mathbb{D})$ that is invariant under $T^{*}_{\phi}-\sum_{i=1}^{k}v_{i}\otimes u_{i}$. Then there exists a $(T^{*}_{\phi}\otimes I_{\mathbb{C}^{p+1}})$-invariant subspace $\mathcal{K}$ of $H^{2}(\mathbb{D},\mathbb{C}^{p+1})$ such that
@@
$\longrightarrow$ 0 as $0\text{ as }n\longrightarrow\infty.$
@@
First, we show that $\mathcal{K}$ is a vector subspace of $H^{2}(\mathbb{D},\mathbb{C}^{p+m})$. Take $(R,H)$ and $(\tilde{R},\tilde{H})$ in $\mathcal{K}$, and a scalar $\alpha\in\mathbb{C}$. Then there exist $F$ and $\tilde{F}\in\mathcal{M}$ such that
@@
(5.9) $F+\alpha\tilde{F}$ $=$ $\mathcal{G}(R+\alpha\tilde{R})+(H+\alpha\tilde{H})$
@@
(5.10) $\mathcal{G}(A_{n}+\alpha\tilde{A_{n}})$ $=$ $\mathcal{G}A_{n}+\alpha\mathcal{G}\tilde{A_{n}}$
@@
$=$ $P_{\mathcal{W}}(T^{*}_{\Phi}P_{\mathcal{M}\ominus\mathcal{W}})^{n}(F+\alpha\tilde{F}).$
@@
and $F,\ R,$ and $H$ satisfy conditions C1, C2, and C3. Then, by the uniqueness of the representation of $F$, we have
@@
Further, we want to emphasize that establishing conditions C1, C2, and C3 for elements of $\mathcal{K}$, once we have the unitary, do not require the assumption that $\mathcal{M}$ is invariant under a perturbation of $T_{\Phi}^{*}$. Indeed, it follows automatically as asserted in Theorem 4.2. Recall that it is a converse of Theorem 4.1. We note that it is a vector-valued analogue of the converse of Theorem 5.1 proved in [DS, Theorem 3.1].
@@
We shall now present the proofs of Theorems 4.3, 4.5 and 4.7. Recall that Theorem 4.3 describes invariant subspaces of some specific finite-rank perturbations of certain Toeplitz operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$. We note that it is a vector-valued extension of a characterization of invariant subspaces of certain finite-rank perturbations of $T_{z}$ on $H^{2}(\mathbb{D})$ given by Das and Sarkar in [DS, Theorems 4.1 & 4.2].
@@
=$\mathcal{M}=\Big{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):(T^{*}_{F_{1}}F,\dots,T^{*}_{F_{p}}F,F)\in\mathcal{N}\Big{\}}.$
@@
It is worth emphasizing that each of these approaches, with their unique mathematical foundations, has sparked new lines of investigation and propelled further advancements in the discipline. Our proof of Theorem 4.5 is more aligned with the techniques used by Das and Sarkar in [DS]. We note that it extends Das and Sarkar’s result to vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$.
@@
[DS, Theorem 5.1] Let $T$ be a bounded operator on a Hilbert space $\mathcal{H}$, and let $\{u_{i}\}_{i=1}^{k}$ and $\{v_{i}\}_{i=1}^{k}$ be two finite sets in $\mathcal{H}$. If a subspace $\mathcal{M}\subset\mathcal{H}$ is invariant under $T-\sum_{i=1}^{k}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T$ with defect at most $k$. Conversely, if $\mathcal{M}$ is almost invariant under $T$ with defect $k$, then $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{k}f_{i}\otimes T^{*}f_{i}$ for every orthonormal basis $\{f_{i}\}_{i=1}^{k}$ of the defect space of $\mathcal{M}$.
@@
Next, we move on to the poof of Theorem 4.7. Recall that Theorem 4.7 describes nearly invariant subspaces of $T_{B,B^{{}^{\prime}}}^{*}$-invariant subspaces of $H^{2}(\mathbb{D},\mathbb{C}^{m}).$ Motivated by the notion of nearly $T^{*}_{z}$-invariant subspaces of $H^{2}(\mathbb{D})$, in [DS], the authors introduced and studied the concept of $T^{*}_{z,B}$-invariant subspaces of $H^{2}(\mathbb{D})$. Clearly, our Theorem 4.7 extends their study to a more general setting in vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. We first recall the definition of a nealy $T^{*}_{\phi,\psi}$-invariant subspace of $H^{2}(\mathbb{D},\mathbb{C}^{m})$ and give a Lemma that records an observation essential for our proof.
@@
This shows that $\mathcal{M}$ is invariant under $T^{*}_{\mathcal{B}}-\sum_{i=1}^{p}T^{*}_{\mathcal{B}}G_{i}\otimes G_{i}$. Therefore, using Theorem 4.1, there exists a $(T^{*}_{B}\otimes I_{\mathbb{C}^{p+m}})$invariant subspace $\mathcal{N}$ of $H^{2}(\mathbb{D},\mathbb{C}^{p+m})$ such that
@@
In [DS], Das and Sarkar describes nearly $T^{*}_{z,B}$-invariant subspaces of $H^{2}(\mathbb{D})$ [DS, Theorem 6.2]. It follows as a corollary to Theorem 4.7 by taking $m=1$ and $B=z$.
6. Proofs of Theorems 4.9 & 4.10, and their general cases
In this section, we shall prove Theorems 4.9 and 4.10 that together characterize nearly $T_{\phi,\psi}^{*}$-invariant subspaces of $H^{2}(\mathbb{D},\mathbb{C}^{m})$ with defect one. We also give the corresponding results for the general case of finite defect $n$.
@@
where $T_{\mathcal{B}}:=T_{B}\otimes I_{\mathbb{C}^{m}},\ T_{\mathcal{B}^{{}^{\prime}}}:=T_{B^{{}^{\prime}}}\otimes I_{\mathbb{C}^{m}}$, and $\mathcal{F}$ is the defect space with dimension $1$ and orthonormal basis $\{J\}$.
@@
$\longrightarrow$ 0 as $0\text{ as }n\longrightarrow\infty.$
@@
(6.7) $\mathcal{K}=\{h:F=BhJ\in\mathcal{M},\ h=f\circ B\ \text{for some}\ f\in H^{2}(\mathbb{D})\ {\rm and}\ ||F||=||h||\big{\}}.$
@@
    =$\mathcal{M}=\Bigg{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=\mathcal{G}R+B\sum_{i=1}^{n}h_{i}J_{i}\ \ {\rm for}\ (R,h_{1},\dots,h_{n})\in\mathcal{K}\Bigg{\}},$
@@
    =$\mathcal{M}=\Bigg{\{}F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=B\sum_{i=1}^{n}h_{i}J_{i}\ \ {\rm for}\ (h_{1},\dots,h_{n})\in\mathcal{K}\Bigg{\}}.$

### v2

@@ Abstract.
In this article, we completely classify invariant subspaces of finite-rank perturbations of a class of Toeplitz operators on vector-valued Hardy spaces. As a consequence, in the vector-valued setting, we characterize invariant and almost invariant subspaces of a class of Toeplitz operators, as well as nearly invariant subspaces associated with certain Blaschke-based operators. We further treat the finite defect case for these nearly invariant subspaces.
@@ 1. Introduction
Perturbations of Toeplitz operators on Hardy spaces have engendered a great deal of interest because–as is well known–they help us understand how small changes in these operators affect their properties, like spectrum or invertibility, which are crucial for solving equations in signal processing and control theory. They’re also key in studying stability in function spaces, where even tiny tweaks can impact convergence or approximation in applications like harmonic analysis or quantum mechanics. The basic ideas in this context are available in the volumes by Böttcher and Silbermann [3] and Rosenblum and Rovnyak [17]. We also refer to the paper by Axler, Chang, and Sarason [1].
@@
In this article, we explore and establish interesting properties and connections between perturbation theory, invariant subspaces, nearly invariant subspaces and almost invariant subspaces in the context of certain Toeplitz operators on vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. See Section 3 for the statements of our main results. It is worth mentioning that Hitt [11] introduced nearly invariant subspaces but called them weakly invariant. Subsequently, Sarason [18] rechristened them with the now universally accepted label nearly invariant subspaces and he introduced new ideas into their study by establishing a new proof via de Branges-Rovnyak spaces. A subspace $\mathcal{M}$ of the Hardy space $H^{2}(\mathbb{D})$ is called nearly invariant under the backward shift operator $T_{z}^{*}$ on $H^{2}(\mathbb{D})$ if $T_{z}^{*}(f)$ belongs to $\mathcal{M}$ whenever $f\in\mathcal{M}$ and vanishes at zero. Since then, numerous generalizations have been proposed in this direction. For further insights into nearly invariant subspaces, one can refer to [4], [5], [8], [6], [9], [10], [12], [14], [15], [16], [18] and the references therein. We also refer to two recent related papers [7] and [13].
@@
A transformative feature of nearly invariant subspaces of the backward shift operator on $H^{2}(\mathbb{D})$ observed and used by Sarason is their connection with a particular rank-one perturbation of the backward shift operator. Inspired by this masterful insight of Sarson, Das and Sarkar [9] have recently investigated the invariant subspaces of certain Toeplitz operators, a generalization of nearly invariant subspaces, and almost invariant subspaces on $H^{2}(\mathbb{D})$ via a classification of invariant subspaces of some finite-rank perturbations of certain Toeplitz operators that they derive in the same paper.
Our work is inspired by the perturbative perspective developed in [9], and we elaborate on this connection in the subsequent sections. In the present paper, we obtain a complete classification of the invariant subspaces of finite-rank perturbations of a class of Toeplitz operators on vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. In doing so, we uncover structural connections between perturbation theory, invariant subspaces, almost invariant subspaces, and nearly invariant subspaces in the vector-valued setting. While our approach is motivated by ideas introduced in [9] for the scalar case, the passage to the vector-valued framework involves additional technical considerations and a more intricate operator-theoretic structure.
The organization of the rest of the paper is as follows. In Section 2, we introduce the notations, and preliminaries used throughout the article. Section 3 contains the statements of our main results.
In Section 4, we prove Theorems 3.1 and 3.2, which provide a classification of invariant subspaces for finite-rank perturbations of a class of Toeplitz operators on the vector-valued Hardy space $H^{2}(\mathbb{D},\mathbb{C}^{m})$. We also establish three consequences of these results:
  1. $(i)$
    a characterization of invariant subspaces of finite-rank perturbations of a class of analytic Toeplitz operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$ (Theorem 3.3);
  2. $(ii)$
    a characterization of almost invariant subspaces of a class of Toeplitz operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$ (Theorem 3.5);
  3. $(iii)$
    a characterization of nearly invariant subspaces associated with certain Blaschke-based operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$ (Theorem 3.7).
Section 5 is devoted to extending Theorem 3.7 to the finite defect case (Theorems 3.9, 3.10, and 5.3). This finite defect analysis does not appear in [9].
Sections 4 and 5 also contain several significant remarks and crucial observations that highlight essential aspects of our findings.
@@ 2. Preliminaries
We shall use $\{E_{i}:1\leq i\leq n\}$ to denote the standard orthonormal basis of $\mathbb{C}^{n}$. Note that each $E_{i}$ can be viewed as a function in $H^{2}(\mathbb{D},\mathbb{C}^{n})$ by identifying it with the constant function that is identically $E_{i}$. Furthermore, $\{E_{j}z^{m}:1\leq j\leq n,\ m\geq 0\}$ forms an orthonormal basis of the Hardy space $H^{2}(\mathbb{D},\mathbb{C}^{n}).$ For $n=1$, we denote $H^{2}(\mathbb{D},\mathbb{C}^{n})$ by $H^{2}(\mathbb{D})$.
@@
The latter identification allows one to view operators between these vector-valued Hardy spaces as row matrices. Suppose $\mathcal{G}$ is a bounded operator from $H^{2}(\mathbb{D},\mathbb{C}^{p})$ to $H^{2}(\mathbb{D},\mathbb{C}^{m})$ and $\mathcal{J}$ is a bounded operator from $H^{2}(\mathbb{D},\mathbb{C}^{r})$ to $H^{2}(\mathbb{D},\mathbb{C}^{m})$, then we shall use the notation $\mathcal{G},\mathcal{J}]$ to denote the bounded operator from $H^{2}(\mathbb{D},\mathbb{C}^{p+r})$ to $H^{2}(\mathbb{D},\mathbb{C}^{m})$ defined by
@@
by a row matrix $G_{1},\dots,G_{m}]$, and call it a matrix with $m$ columns $G_{1},\dots,G_{m}$.
@@
A bounded operator $T$ on a Hilbert space $\mathcal{H}$ is called a $C._{0}$ contraction if for each $h\in\mathcal{H}$, $\|T^{*n}h\|\longrightarrow 0$ as $n\longrightarrow\infty$. A $C._{0}$ isometry is called a pure isometry. We end this section by quoting an interesting result of Benhida and Timotin on $C._{0}$ contraction [2] that is indispensable for our work in the following sections.
@@
[2, Lemma 3.3] Suppose $T$ is a bounded operator on a Hilbert space $\mathcal{H}$. Let $T$ be a $C._{0}$ contraction such that the rank of $I-T^{*}T$ is finite, and let $\mathcal{M}\subset\mathcal{H}$ be a subspace of finite codimension. Then $TP_{\mathcal{M}}$ is a $C._{0}$ contraction, where $P_{\mathcal{M}}$ is the orthogonal projection of $\mathcal{H}$ onto $\mathcal{M}$.
@@ 3. Statements of the main results
The structural insights provided by Theorems 3.1 and 3.2 extend well beyond their immediate context and serve as a foundational tool for analyzing a variety of operator-theoretic phenomena on vector-valued Hardy spaces. We apply these to deduce the following three results.
@@
$\mathcal{M}=\Big\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):(T^{*}_{F_{1}}F,\dots,T^{*}_{F_{p}}F,F)\in\mathcal{N}\Big\},$
@@
Suppose $B$ and $B^{{}^{\prime}}$ are two finite Blaschke products such that $B$ vanishes at zero, and $B$ divides $B^{{}^{\prime}}$ . Let $\mathcal{M}$ be a non-zero nearly $T^{*}_{B,B^{{}^{\prime}}}$-invariant subspace of $H^{2}(\mathbb{D},\mathbb{C}^{m})$ with defect $1$, and let $J$ be a function of unit norm in the defect space.
@@
    $\mathcal{M}=\Big\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=\mathcal{G}R+BhJ\ \ {\rm with}\ \ (R,h)\in\mathcal{K}\Big\},$
@@
    $\mathcal{M}=\Big\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=BhJ\ \ {\rm with}\ \ h\in\mathcal{K}\Big\}.$
@@
    $\mathcal{M}=\Big\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=\mathcal{G}R+BhJ\ \ {\rm for}\ (R,h)\in\mathcal{K}\Big\},$
@@
    $\mathcal{M}=\Big\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=BhJ\ \ {\rm for}\ h\in\mathcal{K}\Big\},$
@@
4. Proofs of Theorems 3.1 & 3.2, and their Applications I, II & III
Before we give the proof of Theorem 3.1, first we state the following result from [9]. It describes invariant subspaces of finite-rank perturbations of a Toeplitz operator $T^{*}_{\phi}$ on $H^{2}(\mathbb{D})$ with the associated symbol $\phi$ an inner function.
@@
[9, Theorem 2.1] Suppose $\phi\in H^{2}(\mathbb{D})$ is an inner function that vanishes at zero; $\{u_{i}\}_{i=1}^{k}$ and $\{v_{i}\}_{i=1}^{k}$ are two orthonormal sets in $H^{2}(\mathbb{D})$. Let $\mathcal{M}$ be a non-zero subspace of $H^{2}(\mathbb{D})$ that is invariant under $T^{*}_{\phi}-\sum_{i=1}^{k}v_{i}\otimes u_{i}$. Then there exists a $(T^{*}_{\phi}\otimes I_{\mathbb{C}^{p+1}})$-invariant subspace $\mathcal{K}$ of $H^{2}(\mathbb{D},\mathbb{C}^{p+1})$ such that
@@
$\longrightarrow$ $0\text{ as }n\longrightarrow\infty.$
@@
[…]
@@
[…]
@@
[…]
@@
and F,R,F,\ R, and $H$ satisfy conditions C1, C2, and C3. Then, by the uniqueness of the representation of $F$, we have
@@
Further, we want to emphasize that establishing conditions C1, C2, and C3 for elements of $\mathcal{K}$, once we have the unitary, do not require the assumption that $\mathcal{M}$ is invariant under a perturbation of $T_{\Phi}^{*}$. Indeed, it follows automatically as asserted in Theorem 3.2. Recall that it is a converse of Theorem 3.1. We note that it is a vector-valued analogue of the converse of Theorem 4.1 proved in [9, Theorem 3.1].
@@
We shall now present the proofs of Theorems 3.3, 3.5 and 3.7. Recall that Theorem 3.3 describes invariant subspaces of some specific finite-rank perturbations of certain Toeplitz operators on $H^{2}(\mathbb{D},\mathbb{C}^{m})$. We note that it is a vector-valued extension of a characterization of invariant subspaces of certain finite-rank perturbations of $T_{z}$ on $H^{2}(\mathbb{D})$ given by Das and Sarkar in [9, Theorems 4.1 & 4.2].
@@
$\mathcal{M}=\Big\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):(T^{*}_{F_{1}}F,\dots,T^{*}_{F_{p}}F,F)\in\mathcal{N}\Big\}.$
@@
It is worth emphasizing that each of these approaches, with their unique mathematical foundations, has sparked new lines of investigation and propelled further advancements in the discipline. Our proof of Theorem 3.5 is more aligned with the techniques used by Das and Sarkar in [9]. We note that it extends Das and Sarkar’s result to vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$.
@@
[9, Theorem 5.1] Let $T$ be a bounded operator on a Hilbert space $\mathcal{H}$, and let $\{u_{i}\}_{i=1}^{k}$ and $\{v_{i}\}_{i=1}^{k}$ be two finite sets in $\mathcal{H}$. If a subspace $\mathcal{M}\subset\mathcal{H}$ is invariant under $T-\sum_{i=1}^{k}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T$ with defect at most $k$. Conversely, if $\mathcal{M}$ is almost invariant under $T$ with defect $k$, then $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{k}f_{i}\otimes T^{*}f_{i}$ for every orthonormal basis $\{f_{i}\}_{i=1}^{k}$ of the defect space of $\mathcal{M}$.
@@
Next, we move on to the poof of Theorem 3.7. Recall that Theorem 3.7 describes nearly invariant subspaces of $T_{B,B^{{}^{\prime}}}^{*}$-invariant subspaces of $H^{2}(\mathbb{D},\mathbb{C}^{m}).$ Motivated by the notion of nearly $T^{*}_{z}$-invariant subspaces of $H^{2}(\mathbb{D})$, in [9], the authors introduced and studied the concept of $T^{*}_{z,B}$-invariant subspaces of $H^{2}(\mathbb{D})$. Clearly, our Theorem 3.7 extends their study to a more general setting in vector-valued Hardy spaces $H^{2}(\mathbb{D},\mathbb{C}^{m})$. We first recall the definition of a nealy $T^{*}_{\phi,\psi}$-invariant subspace of $H^{2}(\mathbb{D},\mathbb{C}^{m})$ and give a Lemma that records an observation essential for our proof.
@@
This shows that $\mathcal{M}$ is invariant under $T^{*}_{\mathcal{B}}-\sum_{i=1}^{p}T^{*}_{\mathcal{B}}G_{i}\otimes G_{i}$. Therefore, using Theorem 3.1, there exists a $(T^{*}_{B}\otimes I_{\mathbb{C}^{p+m}})$ -invariant subspace $\mathcal{N}$ of $H^{2}(\mathbb{D},\mathbb{C}^{p+m})$ such that
@@
In [9], Das and Sarkar describes nearly $T^{*}_{z,B}$-invariant subspaces of $H^{2}(\mathbb{D})$ [9, Theorem 6.2]. It follows as a corollary to Theorem 3.7 by taking $m=1$ and $B=z$.
5. Proofs of Theorems 3.9 & 3.10, and their general cases
In this section, we shall prove Theorems 3.9 and 3.10 that together characterize nearly $T_{\phi,\psi}^{*}$-invariant subspaces of $H^{2}(\mathbb{D},\mathbb{C}^{m})$ with defect one. We also give the corresponding results for the general case of finite defect $n$.
@@
where $T_{\mathcal{B}}:=T_{B}\otimes I_{\mathbb{C}^{m}},\ T_{\mathcal{B}^{{}^{\prime}}}:=T_{B^{{}^{\prime}}}\otimes I_{\mathbb{C}^{m}}\$, and $\mathcal{F}$ is the defect space with dimension $1$ and orthonormal basis $\{J\}$.
@@
$\longrightarrow$ $0\text{ as }n\longrightarrow\infty.$
@@
[…]
@@
    $\mathcal{M}=\Bigg\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=\mathcal{G}R+B\sum_{i=1}^{n}h_{i}J_{i}\ \ {\rm for}\ (R,h_{1},\dots,h_{n})\in\mathcal{K}\Bigg\},$
@@
    $\mathcal{M}=\Bigg\{F\in H^{2}(\mathbb{D},\mathbb{C}^{m}):F=B\sum_{i=1}^{n}h_{i}J_{i}\ \ {\rm for}\ (h_{1},\dots,h_{n})\in\mathcal{K}\Bigg\}.$